### 03 Mar 2026 v2

@@ 4 Algebraic Structure of Categorical Variables
Let $A$ and $B$, be two categorical random variables defined over a common sample space $D$. To fix the idea, let us assume that $A$ stands for income, which we consider to have three possible values, namely $\{l,m,h\}$ for “low”, “middle” and “high”, and that $B$ stands for owning or not some house, which we consider to have two possible values, namely $\{y,n\}$, for “yes” and “no”. It is quite common to produce, out of $A$ and $B$, another random variable $C$, which we will denote as C=ABC=A*B, by the following rule: Given any $p\in D$, then $C(p):=(A(p),B(p))$.
Thus, if $p\in D$ is an individual with middle income, and owns some house, then $C(p)=(m,y)$. Let us observe that the possible values for $C$ is the set
@@
If we denote by $\Sigma_{A}$ the codomain of the random variable $A$ and by $\Sigma_{B}$ the codomain of the random variable $B$, then we get that the codomain of $C$ is $\Sigma_{C}=\Sigma_{A}\times\Sigma_{B}$. Let us also recall that, if $P$ is the probability measure in the sample space $D$, and $P_{A}$, $P_{B}$ are the probability mass function of $A$ and $B$, respectively, then, in the measurable product space $\Sigma_{A}\times\Sigma_{B}$ we have the joint probability mass function $P_{A}\times P_{B}\colon\Sigma_{A}\times\Sigma_{B}\to\mathbb{R}$ given by
@@ 4.1 Formal definitions
Proof.
Consider $A],[B],[C]\in\overline{\cal C}$. For any representatives
@@
For the commutativity, let us observe that $A*B\colon D\to\Sigma_{A*B}=\Sigma_{A}\times\Sigma_{B}$ and $B*A\colon D\to\Sigma_{B*A}=\Sigma_{B}\times\Sigma_{A}$. Hence, we have a bijection between $\Sigma_{A*B}$ and $\Sigma_{B*A}$, and for the partitions we clearly have
${\cal A}\cap{\cal B}={\cal B}\cap{\cal A}.$
Thus, we have [A][B]=[B][A][A]*[B]=[B]*[A]. ∎
@@
Proof.
Let us first observe that any random variable on $D$ having a singleton as the only possible outcome is equivalent to $\Phi$.
Now we observe that $\Sigma_{A*\Phi}=\Sigma_{A}\times\{\phi\}$, so we have a bijection $\Sigma_{A}\to\Sigma_{A*\Phi}$ given by $x\mapsto(x,\phi)$.
Also, the partition corresponding to $\Phi$ is the trivial one ${\cal D}=\{D\}$, and thus the partition of $A*\Phi$ is
${\cal A}\cap{\cal D}={\cal A}.$
Hence, we have $A]*[\Phi]=[A],$ as claimed. ∎
Let us summarize the previous discussion in the following theorem
@@ 4.2 Compatibility of the algebraic and topological structures
$\begin{split}d(X,Y)=1-SU(X,Y)&=1-\frac{2MI(X|I)}{H(X)+X(Y)}\\ &=\frac{H(X)+H(Y)-H(X)+H(X|Y)-H(Y)+H(Y|X)}{H(X)+H(Y)}\\ &=\frac{H(X|Y)+H(Y|X)}{H(X)+H(Y).}\end{split}$ (21)
Also, let us observe that, given categorical random variables $X\colon D\to\Sigma_{X}$ and $Y\colon D\to\Sigma_{Y}$, we have
@@
where $\cal X$} and $\cal Y$} are the partitions of $D$ associated to the variables $A$ and $Y$, respectively.
@@
is continuous with respect to the topology on $\overline{{\cal C}}$} given by d=1SUd=1-SU and the natural product topology induced in $\overline{{\cal C}}\times\overline{{\cal C}}$.
Proof.
Let $X,Y,Z,W$ be categorical random variables on $D$, for simplicity, we will avoid the equivalence class notation, and we will also use the same symbols for their associated partitions. Then we have
@@
Applying item $a)$ from Lemma 2.2, and the inequality given in equation  (12) to the terms in the numerator, we get:
@@
Thus, the mapping is contractive, hence, continuous. ∎
@@ 5 Concluding Remarks
In this paper we have demonstrated that 1SU1-SU, where $SU$ is the entropic correlation between categorical random variables, is a distance metric, thus endowing the space of categorical random variables with a topological structure. We have also showed that a natural joint operation (denoted *) endows the same space with an algebraic structure, concretely, a commutative monoid structure. Finally, we have proven that the topological and algebraic structures are compatible, given that the joint operation is continuous with respect to the topological structure.
The compatibility provides a smooth and straightforward way for the user to interpret and understand the value output from SU. This way, SU has gained intuitivity in the view of the statistics practitioner, giving everyone the ability to operate with and employ this entropic (non-parametric) correlation in much the same way as the Pearson (parametric) correlation has been employed throughout the years.
Equipped with compatible topology and algebraic structure, $SU$ further formalizes CRVs as repositories and proper representations of the concepts conveyed by the variable’s categories. Thus, qualitative variables not only can be counted in their categories, but their mutual distances now tell us about similarities or likely associations. These new possibilities and measures are offered in an environment of mathematical rigor.
The authors are now working to extend the findings to the MSU measure on $n$ categorical variables. Note that the MSU measure could as well be called Multivariable Entropic Correlation, as uncertainties get mutually cancelled when computing correlation.
Acknowledgments
The authors acknowledge Professor Mauricio Poletti for fruitful discussions on the topics of this paper. The authors acknowledge support from FEEI-CONACYT-PROCIENCIA-PRONII and the ARASY project, number ESTR01-23. CES is also affiliated with the Centro de Investigación en Matemática - CIMA, Asunción, Paraguay.

### 30 Mar 2026 v4

@@ 4 Algebraic Structure of Categorical Variables
Let $A$ and $B$ be two categorical random variables defined over a common sample space $D$. To fix the idea, let us assume that $A$ stands for income, which we consider to have three possible values, namely $\{l,m,h\}$ for “low”, “middle” and “high”, and that $B$ stands for owning a house or not, which we consider to have two possible values, namely $\{y,n\}$, for “yes” and “no”. It is quite common to produce, out of $A$ and $B$, another random variable $C$, which we will denote as C=ABC=A*B, by the following rule: Given any $p\in D$, then $C(p):=(A(p),B(p))$.
Thus, if $p\in D$ is an individual with middle income, and owns some house, then $C(p)=(m,y)$. Let us observe that the possible values for $C$ are the set
@@
If we denote by $\Sigma_{A}$ the codomain of the random variable $A$ and by $\Sigma_{B}$ the codomain of the random variable $B$, then we get that the codomain of $C$ is $\Sigma_{C}=\Sigma_{A}\times\Sigma_{B}$. Let us also recall that, if $P$ is the probability measure in the sample space $D$, and $P_{A}$, $P_{B}$ are the probability mass functions of $A$ and $B$, respectively, then, in the measurable product space $\Sigma_{A}\times\Sigma_{B}$ we have the joint probability mass function $P_{A}\times P_{B}\colon\Sigma_{A}\times\Sigma_{B}\to\mathbb{R}$ given by
@@ 4.1 Formal definitions
Proof 4.5.
Consider $A],[B],[C]\in\overline{\cal C}$. For any representatives,
@@
For the commutativity, observe that
$A*B\colon D\to\Sigma_{A*B}=\Sigma_{A}\times\Sigma_{B}\quad\text{and}\quad B*A\colon D\to\Sigma_{B*A}=\Sigma_{B}\times\Sigma_{A}.$
Hence, we have a bijection between $\Sigma_{A*B}$ and $\Sigma_{B*A}$, and for the partitions we clearly have ${\cal A}\cap{\cal B}={\cal B}\cap{\cal A}.$ Thus, we have [A][B]=[B][A][A]*[B]=[B]*[A].
@@
Proof 4.7.
Let us first observe that any random variable on $D$ having a singleton as the only possible outcome is equivalent to $\Phi$. Now, we observe that $\Sigma_{A*\Phi}=\Sigma_{A}\times\{\phi\}$, so we have a bijection $\Sigma_{A}\to\Sigma_{A*\Phi}$ given by $x\mapsto(x,\phi)$. Also, the partition corresponding to $\Phi$ is the trivial one ${\cal D}=\{D\}$, and thus the partition of $A*\Phi$ is ${\cal A}\cap{\cal D}={\cal A}.$ Hence, we have $A]*[\Phi]=[A],$ as claimed.
Let us summarize the previous discussion in the following theorem:
@@ 4.2 Compatibility of the algebraic and topological structures
$\begin{split}d(X,Y)=1-SU(X,Y)&=1-\frac{2MI(X;Y)}{H(X)+H(Y)}\\ &=\frac{H(X)+H(Y)-H(X)+H(X|Y)-H(Y)+H(Y|X)}{H(X)+H(Y)}\\ &=\frac{H(X|Y)+H(Y|X)}{H(X)+H(Y)}\end{split}$ (21)
Also, observe that, given categorical random variables $X\colon D\to\Sigma_{X}$ and $Y\colon D\to\Sigma_{Y}$, we have
@@
where $\cal X$} and $\cal Y$} are the partitions of $D$ associated to the variables $X$ and $Y$, respectively.
@@
is continuous with respect to the topology on $\overline{{\cal C}}$} given by d=1SUd=1-SU and the natural product topology induced on $\overline{{\cal C}}\times\overline{{\cal C}}$.
Proof 4.10.
Let $X,Y,Z,W$ be categorical random variables on $D$, for simplicity, we will avoid the equivalence class notation and we will also use the same symbols for their associated partitions. Then we have
@@
Applying item $a)$ from Lemma 2.2, and the inequality given in equation (12) to the terms in the numerator we get:
@@
Thus, the mapping is contractive, hence, continuous.
@@ 5 Concluding Remarks
In this paper we have demonstrated that 1SU1-SU, where $SU$ is the entropic correlation between categorical random variables, is a distance metric, thus endowing the space of categorical random variables with a topological structure. We have also shown that a natural joint operation (denoted *) endows the same space with an algebraic structure, concretely, a commutative monoid structure. Finally, we have proven that the topological and algebraic structures are compatible, given that the joint operation is continuous with respect to the topological structure.
The compatibility provides a smooth and straightforward way for the user to interpret and understand the output value from SU. This way, SU becomes more intuitive for statistical practitioners, allowing practitioners to work with and employ this entropic (non-parametric) correlation in much the same way as the Pearson (parametric) correlation has been employed over the years.
Equipped with compatible topology and algebraic structure, $SU$ further formalizes categorical random variables (CRVs) as repositories and proper representations of the concepts conveyed by the variable’s categories. Thus, qualitative variables not only can be counted in their categories, but their mutual distances now tell us about similarities or likely associations. These new possibilities and measures are offered in an environment of mathematical rigor.
The authors are now working to extend the findings to the multivariate symmetric uncertainty (MSU) measure on $n$ categorical variables. Note that the MSU measure could as well be called Multivariable Entropic Correlation, as uncertainties get mutually canceled when computing correlation.